Subscribe Subscribe: Network Status
3-27-2015 4:16:46 PM

New UPS equipment is being installed in our Bossier City Data Center on Saturday 3/28/2015 overnight. The maintenance window is planned to begin Saturday 3/28/2015 11:30PM and will last until Sunday 3/29/2015 1:30AM. The maintenance window could end sooner. Minimal downtime is expected. Please check back for updates.

Updates

[2015-03-29 2:15 AM]
The scheduled power maintenance is now complete.

9-26-2013 6:34:12 PM

We are pleased to annouce that all of our web hosting servers are now upgraded to version 5 of our hosting control panel!

Changes and Improvements

  • New Basic Matrix theme to complement the upgrade.
  • RoundCube Webmail upgraded. This fixes recent issues reported for new browsers, adds many feature improvements to the address book, and mail boxes with thousands of messages now load instantly.
  • IPv6 support.
  • Multiple SSL certificates per IP address.
  • Numerous tweaks and bug fixes.
9-12-2012 4:45:33 PM

Maintenance on server vps1002 is scheduled for 11:00 PM CST tonight. We are adding additional disk capacity and rebuilding quotas that are out of sync. ETR: 2 hours.

Updates

2012-09-12 11:00 PM
The maintenance window is now open.

Maintenance Complete

2012-09-12 11:59 PM
Additional disk capacity added, full filesystem error checked, quota resync complete. All quota usage totals look normal again. Server vps1002 is online.

8-2-2012 4:44:29 AM

At approximately 8pm on Wednesday evening August 1st, a reboot of one of our shared hosting servers at the data center triggered a significant system software failure rendering it unbootable. We were aware of the issue almost immediately and worked non-stop to resolve the issue ASAP. Once we repaired the root cause in the file system, the server came back online at 4:01am.

Full service for all customers affected by this outage has resumed as ...

7-14-2012 5:13:26 PM

Our Dallas data center is currently experiencing intermittent connectivity. We are investingating the issue and will post information here as it becomes available.

Updates

2012-07-14 11:55 PM
Intermittent connectivity to select geographic regions continues. Our monitoring indicators show that Basic Matrix equipment appears to be operating correctly at the data center. Once stable connectivity is restored at the data center, we expect all services to resume accordingly.

2012-07-15 2:00 PM
Within 2 hours of the initial ...

6-15-2012 4:19:34 PM

3:23 PM - Our Bossier City data center is experiencing a full outage right now due to a core router failure. Technicians are currently replacing it with a backup router and estimate an ETR of 1 hour. If you are hosted in the Bossier City data center, your website and email are offline, although our servers are operating normally. The data center Internet connection is effectively severed until the core router is replaced.

Since basicmatrix.com is also ...

4-26-2012 5:56:46 PM

The Dallas data center we host in is experiencing a major outage right now, first detected by our monitoring services at 5:10 PM intermittently before becoming a full outage around 5:20 PM. It is unclear at this point if it's network equipment failure or a denial of service (DoS) attack. Our servers are still running, but network connectivity is severed.

Our Bossier City data center is unaffected and completely operational.

As more information is available, we will ...

4-25-2012 7:20:08 PM

Maintenance on the UPS powering all of our equipment in the Dallas data center is scheduled for 11:25 PM CST tonight in response to recurring brownouts experienced lately affecting our uptime. The entire UPS unit will be replaced. ETR: 30 min.

Updates

2012-04-25 10:15 PM
The maintenance window start updated to 5 minutes sooner to allow for shut down time. New Sart: 11:25 PM CST.

2012-04-25 11:25 PM
The maintenance window is now open.

Maintenance Complete

2012-04-25 11:38 PM

4-24-2012 4:40:20 PM

Maintenance on server vps1001 is scheduled for 11:30 PM CST tonight. The quotas need to be resynchronized after a recent power failure at the Dallas data center. ETR: 1 hour.

Updates

2012-04-24 11:30 PM
The maintenance window is now open.

Maintenance Complete

2012-04-25 12:25 AM
Quota resync complete. All quota usage totals look normal again. Server vps1001 is fully online again.

4-11-2012 6:35:52 PM

We are currently experiencing an outage with server vps1001. A runaway process filled it's remaining disk space and now a forced disk scan is checking for errors. We appreciate your patience as we work to get vps1001 back online.

Updates

2012-04-11 7:14 PM
vps1001 is online again. Additional maintenance is required. This maintenance will be performed at midnight tonight.

2012-04-12 1:25 AM
We completed many maintenance tasks , but more still remains, which will be continued ...

3-13-2012 11:57:01 AM

Our Dallas data center experienced brownouts, with downtime first detected by our network monitoring  at 11:18AM. All of our servers in the Dallas data center are online again with the exception of one shared hosting server. It is performing a full disk scan with an approximate completion time of 12:07 PM.

We are currently investigating why the power backup equipment did not protect against the brownouts. Updates will be made to this thread when available.

Update

12-29-2011 10:56:03 AM

One of our primary VPS hosting servers is scheduled today for a memory upgrade, which will effectively double the memory available. The upgrade is scheduled for 5:30 PM CST and the downtime will be less than 15 minutes. During this time, many of our shared hosting customers will experience temporary website or email downtime. We apologize for any inconvenience and appreciate your patience during this necessary upgrade.

Please watch our Facebook, Twitter, and network status pages for ...

12-14-2011 1:23:51 PM
Some of our customers received a high volume of unfiltered spam during a period of approximately 12 hours beginning yesterday afternoon. One of our spam filtering nodes was pushed over capacity and consequently went offline. The node is now back online and spam filtering should be working again for all customers.
8-17-2011 9:22:36 AM

We are experiencing a full phone service outage, and therefore our support and sales lines are currently unavailable.

Please use our other contact options during this time. We will promptly respond to all requests by email.

Updates will be posted to this page.

 

Update: 2011-08-17 10:11 AM CST

Phone service is now restored. Our support and sales lines are open.

5-18-2011 3:10:12 PM

A server in our Dallas data center unexpectedly needed to be rebooted this afternoon. The customers on that server experienced short downtime (less than 10 minutes) and DNS services provided by that server were temporarily disrupted.

We will continue to monitor the server to see if further action is necessary or if this is just a random one time event.

2-14-2011 12:00:00 AM

Some of our clients are currently experiencing difficulty in sending emails to Earthlink customers. While we have not received an official response from EarthLink administrators, we are taking steps today to resolve the immediate issue by reconfiguring our outgoing mail server IP addresses.

We understand this is a serious issue for some customers.  This should resolve the immediate issue as we continue to work with Earthlink's system administrators. Until then, there is a risk that ...

2-2-2011 12:00:00 AM

Issue: At approximately 7:29 AM CST this morning, we experienced an unexpected temporary power outage in our Dallas data center, due to the winter storm passing through Dallas overnight. That data center houses one of our primary DNS servers, which unexpectedly required a disk scan as it rebooted. Disk errors were found and corrected, but the process caused nearly an hour of downtime for that DNS server. Unfortunately, our secondary DNS server in our Shreveport data ...

4-21-2010 12:37:45 PM
Our electricians will be installing surge protectors on the router power circuits at 2:00 AM CST on Friday morning the 23rd. All of the affected routers have dual power feeds that are spread amongst different UPS units and no more than one will be taken down at a time. We do not expect any downtime.
4-16-2010 12:20:12 PM

Starting late in the evening on Thursday April 15th, shared hosting email accounts began experiencing delays both sending and receiving email. This morning we determined the problem was related to a malformed update to the email antivirus definitions. We resolved the issue at approximately 11:20am this morning, April 16th, and email delivery returned to normal. No emails should have been lost during this time as the delayed messages were delivered shortly after the fix was ...

3-3-2010 12:39:53 PM
Dear Shared Hosting Clients,

This morning we experienced a period of downtime beginning at 9:27am and ending just before 10:00am CST. This prime time outage was caused by network provider issue at our data center. Our equipment remained functional and online but experienced a lack of connectivity to the Internet due to a routing issue. We have received the official response below from the data center. We do not anticipate reoccurring network outages from this ...

1-15-2010 2:50:24 PM

In December, 2009, the Optimal Payments Network Operations group started a two-step migration involving their Storage Area Network (SAN), implemented to increase system performance and scalability. The first phase was successfully completed, and the final phase is now scheduled to take place on Sunday, January 17, 2010, beginning at 6:00 a.m. EST and lasting no longer than 120 minutes.

This scheduled downtime will only affect customers using Optimal Payments merchant processing accounts. Clients using our NOVA merchant accounts will ...

1-2-2010 12:58:16 AM

A late bug fix to SpamAssassin has caused it to flag emails received in 2010 as "grossly in the future." This problem has been resolved by the SpamAssassin team, and the update has been pushed to all of our servers. We appologize for the inconvenience, and encourage users of the Learn Ham / Learn Spam / Spam folder sheme to check their Spam folders for legitamate emails that scored too high because of the bug.

Example SpamAssassin Header

10-7-2009 12:21:10 PM

On October 10th at midnight CST, our primary data center in Bedford, TX will be migrating all of our server and networking equipment to their new and improved facility in Dallas, TX. This move is naturally an inconvinence for everyone, but it could not be avoided. The Bedford, TX facility is closing its doors by November 1st.

Basic Matrix sincerely appreciates your patience in this transitional time, and assure you we are doing everything we ...

10-1-2009 12:03:53 PM

On October 1, 2009 a major routing issue caused lack of connectivity to our client servers from 10:25am until 10:50am.  From our initial findings and logs it appears the routing issue created an infinite loop for both inbound and outbound data packets which resulted in lost connectivity. Connectivity to all colocation, dedicated, virtual, and shared hosting servers were affected during this time. No configuration or data was lost during the ~25 minute issue as the servers were operating ...

9-29-2009 12:00:00 AM

A limited outage affecting 3 shared hosting clients on the 66.221.x.x subnet occurred on September 29, 2009. This data center related outage lasted from 2:30pm CST until 5:10pm CST. Investigations as to the cause of this outage are still underway and updates will be posted here when they are available.

During this time no dedicated, virtual, or co-located servers were affected.

9-11-2009 4:02:41 PM

This is a notice for all dedicated, colocation, and shared hosting clients.

The Network Operations group will be performing scheduled maintenance on Saturday, October 10th, 2009 beginning at midnight CST and lasting several hours.

We will be transferring servers to a closed rack at the new data center. During this period expect solid outages lasting in excess of one hour. This is a one-time move and we will be adequately prepared to handle any issues that ...

8-26-2009 2:44:14 PM

At 2pm CST on Wednesday, August 26th a network communications issue required an emergency reboot of a primary server. Disk scans and related recovery routines delayed the system from resuming network services for approximately 30 minutes. Affected shared and dedicated virtual server clients were down during this time.

No permanent data loss was recorded. All systems were back online as of 2:42pm CST.

6-15-2009 8:00:00 AM

There will be a window of downtime between 5:30pm - 6:30pm lasting long enough to perform a high priority disk integrity scan of the shared hosting disk space.

This interruption will be kept to a minimum. If you are experiencing downtime after 6:30pm, check here for updates and report such outages.

****************************** UPDATE ******************************
All file allocation table problems were repaired. The downtime affected only our shared hosting server. System was brought back online by 6:10pm CST.

6-10-2009 7:54:05 PM

At 7:20pm CST on 6-10-2009, an unexpected power inconsistency caused a selection of our virtual servers to shut down. We are working to remedy the situation and scanning for any possible data corruption.

We will post more information here as it becomes available.

********************* UPDATE ************************
All virtual servers were brought back online by 8pm CST. We are running scans to detect and repair any data corruption caused by the outage.

This service issue has been ...

6-1-2009 8:52:16 AM

We are experiencing a surge of spam emails being sent from spam networks to our servers. Our usual spam filters are in place preventing this spam from being delivered to mailboxes but the sheer volume combined with processing time are causing delayed email delivery. No e-mails are lost, only temporarily delayed.

Customers using shared hosting services may experience email delivery delays varying from several minutes to several hours as our technicians work to resolve this ...

3-12-2009 11:00:51 AM

Maintenance is scheduled for all shared hosting accounts on Thursday, March 23, 2009 between 11:00pm-12:00am. Service interruption may be experienced for up to 30min within the maintenance window.

10-29-2008 4:50:13 PM

This is a notice for all credit card processing customer using integrated merchant accounts.

The Network Operations group will be performing scheduled maintenance on Sunday, November 2nd, 2008, beginning at 5:00 a.m. EST and lasting no longer than 3 hours.

During this period, you will be unable to process online transactions, nor will you be able to access the merchant admin interface.

Please accept our apologies for any inconvenience this may cause you. If you have any ...

10-8-2008 1:53:30 PM

All customers have been migrated to new servers with significantly more processing and diskspace capacity.